OpenStudy (anonymous):

3ln2 + 2ln4 write each as a single logarithm

OpenStudy (anonymous):

Properties of logarithms!!! \[\huge \ln x^p=p\ln x\]

OpenStudy (anonymous):

I know the first thing I'm supposed to do is ln2^3+ln4^2 but then aren't you supposed to multiply them?

OpenStudy (abb0t):

\(aln(b) = \ln(b)^a\) also, \(\ln(a)-\ln(b) = \ln \frac{ a }{ b }\) \(\ln(a)+\ln(b) = \ln(ab)\)

OpenStudy (anonymous):

Later :) \[\huge \ln2^3+\ln4^2=\ln8+\ln16\]And now use this property \[\huge \ln(ab)=\ln(a)+\ln(b)\]

OpenStudy (anonymous):

So is the answer 128?

OpenStudy (anonymous):

No. Logarithms don't just suddenly disappear, silly ^.^

OpenStudy (anonymous):

So how do I get the final answer?1

OpenStudy (anonymous):

Well, using that property \[\huge \ln(8)+\ln(16) = \ln(8\cdot 16)\]

OpenStudy (anonymous):

I just did that! So \[\ln 128\]

OpenStudy (anonymous):

Yeah, it's ln(128) not just 128 :P

OpenStudy (anonymous):

why does it say in my book that the answer is 7ln2 then?

OpenStudy (anonymous):

Well, maybe you could use the fact that \[\huge 128 = 2^7\] and the first property I mentioned ^.^

OpenStudy (anonymous):

Well that makes sense

OpenStudy (anonymous):

I do love making sense ^.^

OpenStudy (anonymous):

Another question I have is, 3ln5+4lnx

OpenStudy (anonymous):

I solved it, I just want to know if I got the correct answer

OpenStudy (anonymous):

Go for it :)

OpenStudy (anonymous):

Okay I got \[\ln 125x ^{4}\]

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Ah thank you!